# Mamba Can Learn Low-Dimensional Targets In-Context via Test-Time Feature Learning

### v2

@@ Appendix B One-Step Gradient Descent on the Mamba Layer
${\bm{\gamma}}^{*}$ $=\frac{2\eta}{T_{1}}\sum_{t\in[T_{1}]}\left[\Big(y^{t}-f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)\Big)\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)\right]$
$=\frac{2\eta}{T_{1}}\sum_{t\in[T_{1}]}y^{t}\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)$
$\quad-\frac{2\eta}{T_{1}}\sum_{t\in[T_{1}]}f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)\Big)\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right).$
@@
$\quad\left|f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)\right|$
$={\mathrm{ReLU}}\left(N_{\mathrm{pt}}^{-1}\sum_{j=1}^{N_{\mathrm{pt}}}G_{j,n+1}({\bm{Z}})y_{j}^{t}\phi\left({\bm{x}}_{j}^{t}\right)^{\top}\left({\bm{\gamma}}(0)\odot\phi\left({\bm{x}}^{t}\right)\right)\right)$
@@
$\quad\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)$
$=\mathbbm{1}\left[\sum_{j=1}^{N_{\mathrm{pt}}}G_{j,N_{\mathrm{pt}}+1}\left({\bm{Z}}^{t}\right)y_{j}^{t}\phi\left({\bm{x}}_{j}^{t}\right)^{\top}\left({\bm{\gamma}}(0)\odot\phi\left({\bm{x}}^{t}\right)\right)>0\right]$
@@
$\left\|\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)\right\|$ $\leq N_{\mathrm{pt}}^{-1}\sum_{h=1}^{N_{\mathrm{pt}}}G_{j,n+1}({\bm{Z}})\left|y_{j}^{t}\right|\left\|\phi\left({\bm{x}}_{j}^{t}\right)\right\|\left\|\phi\left({\bm{x}}^{t}\right)\right\|$
@@
${\bm{\gamma}}^{*}=\frac{2\eta}{T_{1}}\sum_{t\in[T_{1}]}y^{t}\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)+{\tilde{\mathcal{O}}}\left(\eta d^{-2C_{b}+4}\right).$
@@ B.1 Estimation of Label-Gradient Correlation
$\quad N_{\mathrm{pt}}^{-1}\sum_{j=1}^{N_{\mathrm{pt}}}G_{j,N_{\mathrm{pt}}+1}\left({\bm{Z}}^{t}\right)y_{j}^{t}\phi\left({\bm{x}}_{j}^{t}\right)^{\top}\left({\bm{\gamma}}(0)\odot\phi\left({\bm{x}}^{t}\right)\right)$
@@
$\quad N_{\mathrm{pt}}^{-1}\sum_{j=1}^{N_{\mathrm{pt}}}G_{j,N_{\mathrm{pt}}+1}\left({\bm{Z}}^{t}\right)y_{j}^{t}\phi\left({\bm{x}}_{j}^{t}\right)^{\top}\left({\bm{\gamma}}(0)\odot\phi\left({\bm{x}}^{t}\right)\right)$
@@
$\quad+{\tilde{\mathcal{O}}}\left(d^{-C_{b}+2}N_{\mathrm{pt}}^{-1/2}\right).$
@@ B.2 Characterization of Updated Parameter
$\quad\left\|\frac{1}{T_{1}}\sum_{t=1}^{T_{1}}y^{t}\mathbbm{1}\left[a_{0}\gamma^{2}+a_{1}\left\langle{\bm{\beta}}^{t},{\bm{x}}^{t}\right\rangle+a_{2}\gamma\mathrm{He}_{2}\left(\left\langle{\bm{\beta}}^{t},{\bm{x}}^{t}\right\rangle\right)\right]\phi\left({\bm{x}}^{t}\right)\odot\psi\left({\bm{\beta}}^{t},a_{0},a_{1},a_{2}\right)-{\bm{c}}\right\|$
$={\tilde{\mathcal{O}}}\left(d^{-C_{b}+1}T_{1}^{-1/2}\right),$
@@
$\quad\frac{1}{T_{1}}\sum_{t=1}^{T_{1}}y^{t}\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}(0),{\bm{u}}(0),{\bm{v}}(0),{\bm{a}}(0)\right)$
$=\mathbb{E}_{{\bm{\beta}}\sim\mathrm{Unif}(S_{r})}\left[\psi\left({\bm{\beta}},a_{0},a_{1},a_{2}\right)\odot\psi\left({\bm{\beta}},b_{0},b_{1},b_{2}\right)\right]$
$\quad+{\tilde{\mathcal{O}}}\left(d^{-C_{b}+2}N_{\mathrm{pt}}^{-1/2}\right)+{\tilde{\mathcal{O}}}\left(d^{-C_{b}+1}T_{1}^{-1/2}\right),$

### v3

@@ Appendix B One-Step Gradient Descent on the Mamba Layer
${\bm{\gamma}}^{*}$ $=\frac{2\eta}{T_{1}}\sum_{t\in[T_{1}]}\left[\Big(y^{t}-f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)\Big)\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)\right]$
$=\frac{2\eta}{T_{1}}\sum_{t\in[T_{1}]}y^{t}\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)$
$\quad-\frac{2\eta}{T_{1}}\sum_{t\in[T_{1}]}f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)\Big)\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right).$
@@
$\quad\left|f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)\right|$
$={\mathrm{ReLU}}\left(N_{\mathrm{pt}}^{-1}\sum_{j=1}^{N_{\mathrm{pt}}}G_{j,n+1}({\bm{Z}})y_{j}^{t}\phi\left({\bm{x}}_{j}^{t}\right)^{\top}\left({\bm{\gamma}}^{(0)}\odot\phi\left({\bm{x}}^{t}\right)\right)\right)$
@@
$\quad\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)$
$=\mathbbm{1}\left[\sum_{j=1}^{N_{\mathrm{pt}}}G_{j,N_{\mathrm{pt}}+1}\left({\bm{Z}}^{t}\right)y_{j}^{t}\phi\left({\bm{x}}_{j}^{t}\right)^{\top}\left({\bm{\gamma}}^{(0)}\odot\phi\left({\bm{x}}^{t}\right)\right)>0\right]$
@@
$\left\|\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)\right\|$ $\leq N_{\mathrm{pt}}^{-1}\sum_{h=1}^{N_{\mathrm{pt}}}G_{j,n+1}({\bm{Z}})\left|y_{j}^{t}\right|\left\|\phi\left({\bm{x}}_{j}^{t}\right)\right\|\left\|\phi\left({\bm{x}}^{t}\right)\right\|$
@@
${\bm{\gamma}}^{*}=\frac{2\eta}{T_{1}}\sum_{t\in[T_{1}]}y^{t}\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)+{\tilde{\mathcal{O}}}\left(\eta d^{-2C_{b}+4}\right).$
@@ B.1 Estimation of Label-Gradient Correlation
$\quad N_{\mathrm{pt}}^{-1}\sum_{j=1}^{N_{\mathrm{pt}}}G_{j,N_{\mathrm{pt}}+1}\left({\bm{Z}}^{t}\right)y_{j}^{t}\phi\left({\bm{x}}_{j}^{t}\right)^{\top}\left({\bm{\gamma}}^{(0)}\odot\phi\left({\bm{x}}^{t}\right)\right)$
@@
$\quad N_{\mathrm{pt}}^{-1}\sum_{j=1}^{N_{\mathrm{pt}}}G_{j,N_{\mathrm{pt}}+1}\left({\bm{Z}}^{t}\right)y_{j}^{t}\phi\left({\bm{x}}_{j}^{t}\right)^{\top}\left({\bm{\gamma}}^{(0)}\odot\phi\left({\bm{x}}^{t}\right)\right)$
@@
$=y^{t}\mathbbm{1}\left[a_{0}\gamma^{2}+a_{1}\left\langle{\bm{\beta}}^{t},{\bm{x}}^{t}\right\rangle+a_{2}\gamma\mathrm{He}_{2}\left(\left\langle{\bm{\beta}}^{t},{\bm{x}}^{t}\right\rangle\right)>0\right]\phi({\bm{x}}^{t})\odot\psi\left({\bm{\beta}}^{t},a_{0},a_{1},a_{2}\right)+{\tilde{\mathcal{O}}}\left(d^{-C_{b}+2}N_{\mathrm{pt}}^{-1/2}\right).$
@@ B.2 Characterization of Updated Parameter
$\left\|\frac{1}{T_{1}}\sum_{t=1}^{T_{1}}y^{t}\mathbbm{1}\left[a_{0}\gamma^{2}+a_{1}\left\langle{\bm{\beta}}^{t},{\bm{x}}^{t}\right\rangle+a_{2}\gamma\mathrm{He}_{2}\left(\left\langle{\bm{\beta}}^{t},{\bm{x}}^{t}\right\rangle\right)\right]\phi\left({\bm{x}}^{t}\right)\odot\psi\left({\bm{\beta}}^{t},a_{0},a_{1},a_{2}\right)-{\bm{c}}\right\|={\tilde{\mathcal{O}}}\left(d^{-C_{b}+1}T_{1}^{-1/2}\right),$
@@
$\quad\frac{1}{T_{1}}\sum_{t=1}^{T_{1}}y^{t}\nabla_{\bm{\gamma}}f\left({\bm{Z}}^{t};{\bm{\gamma}}^{(0)},{\bm{u}}^{(0)},{\bm{v}}^{(0)},{\bm{a}}^{(0)}\right)$
$=\mathbb{E}_{{\bm{\beta}}\sim\mathrm{Unif}(S_{r})}\left[\psi\left({\bm{\beta}},a_{0},a_{1},a_{2}\right)\odot\psi\left({\bm{\beta}},b_{0},b_{1},b_{2}\right)\right]+{\tilde{\mathcal{O}}}\left(d^{-C_{b}+2}N_{\mathrm{pt}}^{-1/2}\right)+{\tilde{\mathcal{O}}}\left(d^{-C_{b}+1}T_{1}^{-1/2}\right),$